To compare the effect of stress caused by noise on the ability to perform a simple task, 78 subjects were divided into two groups—34 subjects as a control group, and 44 subjects as the experimental group. Although each subject performed the task, the experimental group had to perform the task while loud rock music was played. The time to finish the task was recorded for each subject and the following summary was obtained.

Control Experimental
n 34 44
x 18 minutes 25 minutes
s 5 minutes 11 minutes

(a)

Find a 99% confidence interval for the difference in mean completion times (in minutes) for these two groups. (Calculate the difference as the control group − the experimental group. Round your answers to two decimal places.)
Answer #1

a)

for 99 % CI value of z= 2.58
margin of error E=z*std error = 4.809
lower bound=(x1-x2)-E = -11.82
Upper bound=(x1-x2)+E = -2.18
from above 99% confidence interval for population mean =(-11.82 , -2.18)